# Code-switching Speech Recognition Under the Lens: Model- and Data-Centric Perspectives

### v1

@@ Abstract
Code-switching automatic speech recognition (CS-ASR) presents unique challenges due to language confusion introduced by spontaneous intra-sentence switching and accent bias that blurs the phonetic boundaries. Although the constituent languages may be individually high-resource, the scarcity of annotated code-switching data further compounds these challenges. In this paper, we systematically analyze CS-ASR from both model-centric and data-centric perspectives. By comparing state-of-the-art algorithmic methods, including language-specific processing and auxiliary language-aware multi-task learning, we discuss their varying effectiveness across datasets with different linguistic characteristics. On the data side, we first investigate TTS as a data augmentation method. By varying the textual characteristics and speaker accents, we analyze the impact of language confusion and accent bias on CS-ASR. To further mitigate data scarcity and enhance textual diversity, we propose a prompting strategy by simplifying the equivalence constraint theory (SECT) to guide large language models (LLMs) in generating linguistically valid code-switching text. The proposed SECT outperforms existing methods in ASR performance and linguistic quality assessments, generating code-switching text that more closely resembles real-world code-switching text. When used to generate speech–text pairs via TTS, SECT proves effective in improving CS-ASR performance. Our analysis of both model- and data-centric methods underscores that effective CS-ASR requires strategies to be carefully aligned with the specific linguistic characteristics of the code-switching data.
@@ I Introduction
Code-switching involves the alternation between two or more languages within spontaneous multilingual speech [1]. Intra-sentence code-switching occurs when the language changes within a single sentence, while inter-sentential code-switching involves language alternation at sentence boundaries and is commonly treated as multilingual [2, 3]. Despite significant advancements in monolingual speech processing tasks, such as automatic speech recognition (ASR) and text-to-speech synthesis (TTS) [4, 5, 6, 7], code-switching speech processing remains challenging due to both the complex linguistic characteristics inherent to code-switched speech and the scarcity of annotated data.
@@
The inherent challenges in code-switching speech processing can be broadly attributed to two primary factors, which we define as accent bias and language confusion. Speakers may exhibit native-like fluency in multiple languages and seamlessly alternate between them during spontaneous conversation. However, accent, particularly common among L2 speakers, can blur the phonetic boundaries between the two languages. We define this phenomenon as accent bias, which complicates both acoustic modeling and the mapping from acoustic features to token embeddings. In the meantime, code-switching introduces language confusion within the matrix language, the dominant or structural language in a code-switching utterance. This confusion is primarily associated with text and arises from the insertion of elements from an embedded language that disrupts the expected syntactic and phonological patterns during language modeling [8]. Moreover, in naturally bilingual communities, the matrix and embedded languages may alternate, resulting in increasingly ambiguous boundaries between them. This linguistic complexity presents additional challenges. As a result, even large-scale pre-trained multilingual speech models, such as Whisper [9], exhibit degraded performance on code-switching data, despite demonstrating high performance in both the matrix and embedded languages individually.
@@
Accent bias has typically been mitigated from a data-centric perspective through adaptation with accented speech [10, 11]. Existing studies have attempted to address the language confusion in code-switching speech recognition (CS-ASR) by integrating language discriminative information. One line of research has sought to factorize the recognition process into separate monolingual components [12, 13]. To effectively coordinate the language-specific processing, an additional divide-and-conquer strategy, such as the mixture-of-experts (MoE) framework or factorization methods, is often required [12, 13, 14, 15, 16]. Another important stream of work has pursued a unified multi-task paradigm [17, 18, 19, 20, 21]. Such approaches involve multi-tasking learning with an auxiliary language identification (LID) or diarization (LD) module to enrich the unified model with frame- or token-level language information [18, 22, 3].
@@
Prior research has explored data augmentation strategies to mitigate data scarcity for low-resource and code-switching ASR [23, 24, 25]. Traditional data augmentation methods for ASR include speed perturbation and SpecAugment [26, 27]. These approaches expand training data by altering the acoustic properties of the original speech signal, such as modifying the speaking rate or applying spectrogram masking. Voice conversion (VC) and TTS provide another solution by introducing speaker variability or generating speech signals from text data. Pre-trained VC and TTS systems typically underperform for low-resource languages due to limited tokenizer exposure and insufficient linguistic coverage [7]. In contrast, code-switching scenarios involving high-resource language pairs, such as English and Mandarin, moderately circumvent this limitation, as both languages are well-represented in large-scale TTS training corpora [28, 29]. Consequently, synthesizing code-switching speech using TTS and VC systems pre-trained on high-resource constituent languages offers a practical and effective data augmentation strategy to improve the performance of CS-ASR [30]. Existing research has also explored synthesizing code-switching text to increase textual diversity for TTS-based data augmentation in ASR training [31, 32]. Earlier approaches typically relied on rule-based frameworks or generative models such as generative adversarial networks (GANs) [32, 31]. More recently, LLMs have been adopted for this task [33, 30] and have proven effective in generating code-switching text data that more aligns with real-world code-switching sentences.
In this study, we investigate CS-ASR from both model- and data-centric perspectives. On the model side, we conduct a comprehensive review and comparison of recent algorithmic approaches, including language-specific processing and multi-task learning with auxiliary language-aware objectives. On the data side, TTS-based data augmentation is first explored to generate paired speech–text code-switching data. We next disentangle the effects of language confusion and accent bias and analyze their impact on CS-ASR by synthesizing speech with varied textual characteristics and accents. To enhance textual diversity, we introduce a prompting strategy that simplifies the equivalence constraint theory (SECT) to guide LLMs in generating linguistically valid code-switching text [34]. The generated text can be further leveraged with TTS to create diverse and effective training data for CS-ASR. By comparing model- and data-centric approaches, we summarize practical strategies for different code-switching scenarios, offering insight for advancing CS-ASR research and applications.
@@ II-A Unified model and auxiliary tasks
However, the above approach rarely makes explicit use of language information. To mitigate this limitation, recent studies have explored enriching the model via multi-task learning with auxiliary language-aware tasks, such as LD or LID [17, 18, 19]. These tasks are typically performed at the frame or token level to achieve fine-grained language change detection. The CS-ASR model is then optimized with a multi-task objective, formulated as the weighted sum of the ASR loss and the auxiliary language-aware loss as
@@ II-B Language-specific Modules
Another important line of research focuses on incorporating language-specific processing within CS-ASR models. Language-specific encoders process the acoustic feature vectors $\mathbf{X}$ separately into hidden representations $\mathbf{H}=(\mathbf{h}_{t}\in\mathbb{R}^{D}\mid t=1,\ldots,T^{\prime})$ before performing a weighted sum of them. The weights $\mathbf{A}^{mat/emb}=(a_{t}\in[0,1]\mid t=1,\ldots,T)$ are frame-level language posteriors computed by a built-in LID module. These are computed via
@@
The auxiliary language-aware task is typically assigned a relatively small weight (i.e., high $\alpha$ value) compared to the ASR task during training, indicating its minor role in guiding the optimization. By contrast, language-specific modules contribute symmetrically across languages. Each module processes features in the same manner [12, 15], regardless of the relative proportions of matrix and embedded languages in the training data. In the training stage, language-specific modules may require pre-training on monolingual data to learn language-discriminative information before being applied to code-switching data, whereas multi-task learning can be performed solely on code-switching data [18, 19, 21]. During inference, the auxiliary language branch is typically inactive, the decoding process is thus identical to that of a unified model. In contrast, language-specific modules require two passes through the encoder or decoder, introducing additional computational overhead and higher decoding latency. For example, the bi-encoder method exhibits real-time factors (RTF) and latency more than twice those of a model with a single encoder. These differences highlight that auxiliary tasks can enhance CS-ASR with minimal overhead, while language-specific modules deliver more explicit language discrimination at the expense of efficiency.
@@ III-A Data Scarcity and Monolingual Utilization
To address this challenge, existing research has leveraged resources from the constituent languages and explored methods for synthesizing code-switching text and speech. Several studies have shown that monolingual data, particularly from the matrix language, can support the development of CS-ASR systems [12, 13, 36, 37, 38, 39]. For example, [36, 37, 38] investigated building CS-ASR models using only monolingual data. These approaches typically employ language-specific connectionist temporal classification (CTC) [40, 37], where decoding is first performed independently for each language and then combined through joint decoding. In this process, tokens in the matrix-language sequence are substituted with the corresponding embedded-language tokens when appropriate. However, these works also indicate that subsequent fine-tuning on code-switching data provides substantial performance gains. Moreover, an excessive reliance on monolingual data can result in training imbalance and degrade performance on code-switching speech. Therefore, augmenting code-switching data is crucial for effectively addressing this scarcity.
@@ III-B Code-switching speech synthesis
Figure 2: The CosyVoice2 model used for TTS; only the LLM module within CosyVoice2 is updated during fine-tuning.
@@
To analyze the effects of text style and accent bias on CS-ASR, we simulate diverse code-switching speech–text pairs via TTS. The synthetic speech is generated using varying combinations of target and reference text and speech domains, such that it inherits the acoustic characteristics and speaker identity from the reference (i.e., prompt) speech while adhering to the textual pattern of the target text. This design allows us to isolate and evaluate the individual and combined effects of stylistic and accentual variation on ASR robustness.
@@ III-C Code-switching text generation
CS-ASR is inherently a cross-lingual task, as it involves dynamic alternation between two or more languages within a single utterance or conversation. A natural solution to this challenge is to develop a multilingual ASR model equipped with text translation capabilities across language pairs. However, such an approach typically demands large-scale supervised multilingual and parallel data, which are often unavailable, particularly for spontaneous speech commonly observed in code-switching scenarios. Consequently, the applicability of translation-augmented multilingual ASR models remains limited in real-world settings due to severe data scarcity.
An alternative strategy focuses on directly optimizing ASR models with code-switching data. However, this approach is often limited by data scarcity. Although TTS can alleviate the shortage of speech resources, generating diverse and natural code-switching text remains a major challenge. This difficulty arises from the wide variability in code-switching patterns, which may involve arbitrary combinations of syntactic structures, lexical substitutions, and language change points. Consequently, synthesizing high-quality code-switching text is crucial, as it forms the basis for constructing paired speech–text data via TTS.
@@
LLMs present a promising solution for this task due to their inherent multilingual and translation capabilities. However, unconstrained generation may lead to unnatural or linguistically implausible switches. To address this, linguistic constraints such as the ECT can be incorporated to guide generation [34, 44, 31, 33, 45]. ECT constrains code-switching to occur only at points where the grammatical structures of both languages align. When implemented in prompt-based generation, ECT can be simplified since LLMs already internalize linguistic regularities to maintain syntactic compatibility and naturalness.
@@
As depicted in Figure 3, monolingual sentences are translated from a high-resource matrix language (Mandarin) into their English-Mandarin code-switching counterparts via the prompt-guided LLM. This process is governed by SECT, which is formulated as three rules. The first rule specifies the matrix language and prevents the LLM from making substantial modifications to the original sentence. The second rule introduces a constraint to control the degree of code mixing, which can be adjusted depending on the desired characteristics of the target code-switching text. Finally, the third rule ensures that the generated code-switching sentences adhere to the syntactic structure of the matrix language while embedding content words from the embedded language at linguistically appropriate switch points. To further guide the model, we adopt an in-context learning strategy by providing three illustrative examples within the prompt. Each example consists of an input and output, being a monolingual sentence and its code-switching counterpart.
Compared with existing LLM-based approaches for code-switching text generation [33, 30], SECT offers a more convenient solution by exploiting the LLM’s intrinsic knowledge and eliminating the need for feeding parallel matrix–embedded language text into each prompt. By leveraging the generative capability of LLMs under SECT constraints, the proposed method synthesizes high-quality code-switching text directly from monolingual data, which can then be used for downstream ASR training.
@@ IV-A Datasets
The two code-switching datasets differ primarily in accent and syntactic structure. Since the ASRU dataset is recorded in mainland China, data within this corpus is characterized by a Chinese accent and contains Mandarin-centric text, with a grammatical structure primarily framed in Mandarin and embedded English segments. Sentences within the ASRU training and development sets, on average, contains 1.6 English words and 8.6 Chinese characters. As opposed to the ASRU data, the SEAME dataset is recorded in Singapore and Malaysia and features speakers with South-East Asian accents. We also compute the code mixing index (CMI) as defined in [49], to measure the degree of language mixing of code-switching text data. The CMI values of the ASRU and SEAME training sets are 17.0 and 13.6, respectively, while the CMI of SEAME increases to 24.2 when monolingual utterances are excluded. Therefore, SEAME data contains more frequent code-switching than the ASRU dataset, largely due to the bilingual education systems and language policies in these regions [50]. The higher frequency and fluidity of language switching suggest that SEAME poses greater challenges for code-switching ASR, particularly as the matrix and embedded languages can alternate across sentences [51, 52].

### v2

@@ Abstract
Code-switching automatic speech recognition (CS-ASR) presents unique challenges due to language confusion that violates the monolingual assumption and accent bias that blurs the phonetic boundaries. Although the constituent languages may be individually high-resource, the scarcity of annotated code-switching data further compounds these challenges. In this paper, we systematically analyze CS-ASR from both model-centric and data-centric perspectives. By comparing state-of-the-art algorithmic methods, including language-specific processing and auxiliary language-aware multi-task learning, we discuss their varying effectiveness across datasets with different linguistic characteristics. On the data side, we first investigate TTS as a data augmentation method. By varying the textual characteristics and speaker accents, we analyze the impact of language confusion and accent bias on CS-ASR. To further mitigate data scarcity and enhance textual diversity, we propose a prompting strategy by simplifying the equivalence constraint theory (SECT) to guide large language models (LLMs) in generating linguistically valid code-switching text. The proposed SECT outperforms existing methods in ASR performance and linguistic quality assessments, generating code-switching text that more closely resembles real-world code-switching text. When used to generate speech–text pairs via TTS, SECT proves effective in improving CS-ASR performance. Our analysis of both model- and data-centric methods underscores that effective CS-ASR requires strategies to be carefully aligned with the specific linguistic characteristics of the code-switching data.
@@ I Introduction
Code-switching involves the alternation between two or more languages within spontaneous multilingual speech [47]. Intra-sentence code-switching occurs when the language changes within a single sentence, while inter-sentential code-switching involves language alternation at sentence boundaries and is commonly treated as multilingual [42, 26]. Despite significant advancements in monolingual speech processing tasks, such as automatic speech recognition (ASR) and text-to-speech synthesis (TTS) [14, 40, 44, 11], code-switching speech processing remains challenging. The challenges arise from both the complex linguistic characteristics inherent to code-switched speech and the scarcity of annotated data.
@@
The inherent challenges in code-switching speech processing can be broadly attributed to two primary factors. We define the inherent challenges as accent bias and language confusion as depicted in Figure 1. Speakers may exhibit native-like fluency in multiple languages and seamlessly alternate between them during spontaneous conversation. However, accent, particularly common among L2 speakers, can blur the phonetic boundaries between the two languages. We define this phenomenon as accent bias, which complicates both acoustic modeling and the mapping from acoustic features to token embeddings. In parallel, code-switching introduces language confusion within the matrix language, the dominant or structural language in a code-switching utterance. This confusion is primarily associated with text. It arises from the insertion of elements from an embedded language that disrupts the expected syntactic and phonological patterns during language modeling [15]. Moreover, in naturally bilingual communities, the matrix and embedded languages may alternate. This alternation leads to increasingly ambiguous boundaries between the two languages. This linguistic complexity presents additional challenges. As a result, even large-scale pre-trained multilingual speech models, such as Whisper [43], exhibit degraded performance on code-switching data. This performance drop occurs despite their high performance in both the matrix and embedded languages individually.
Figure 1: Overview of existing methods in code-switching ASR.
@@
Accent bias has typically been mitigated from a data-centric perspective through adaptation with accented speech [49, 61]. Existing studies have attempted to address the language confusion in code-switching speech recognition (CS-ASR) by integrating language discriminative information. One line of research has sought to factorize the recognition process into separate monolingual components [29, 58]. To effectively coordinate the language-specific processing, an additional divide-and-conquer strategy is often required. Common choices include mixture-of-experts (MoE) framework or factorization methods [29, 58, 50, 65, 48, 53]. Another important stream of work has pursued a unified multi-task paradigm [64, 51, 27, 25, 28]. Such approaches involve multi-tasking learning with an auxiliary language identification (LID) or diarization (LD) module to enrich the unified model with frame- or token-level language information [51, 22, 26]. In addition, post-processing techniques have proven effective in mitigating language confusion. These include late fusion with a language model (LM) and error correction [8, 28].
@@
Prior research has explored data augmentation strategies to mitigate data scarcity for low-resource and code-switching ASR [4, 59, 35]. Traditional data augmentation methods for ASR include speed perturbation and SpecAugment [19, 37]. These approaches expand training data by altering the acoustic properties of the original speech signal, such as modifying the speaking rate or applying spectrogram masking. Voice conversion (VC) and TTS provide another solution by introducing speaker variability or generating speech signals from text data. Pre-trained VC and TTS systems typically underperform for low-resource languages due to limited tokenizer exposure and insufficient linguistic coverage [11]. In contrast, code-switching scenarios involving high-resource language pairs, such as English and Mandarin, moderately circumvent this limitation, as both languages are well-represented in large-scale TTS training corpora [66, 54]. Consequently, synthesizing code-switching speech using TTS and VC systems pre-trained on high-resource constituent languages offers a practical and effective data augmentation strategy to improve the CS-ASR performance  [56]. Existing research has also explored synthesizing code-switching text to increase textual diversity for TTS-based data augmentation in ASR training [7, 63, 17]. Earlier approaches typically relied on rule-based frameworks or generative models such as generative adversarial networks (GANs) [63, 7]. More recently, large language models (LLMs) have been adopted for this task [20, 17, 56]. They have proven effective in generating code-switching text data that more aligns with real-world code-switching sentences.
In this paper, we investigate CS-ASR from both model- and data-centric perspectives. The contributions of this work are summarized as follows:
  • Model-centric analysis. We systematically review and compare recent algorithmic approaches, including language-specific processing and multi-task learning with language-aware tasks. While our analysis reveals the effectiveness of algorithmic methods, their performance is often limited by accent bias. This observation motivates a data-centric investigation into the respective impacts of accent bias and language confusion.
  • Data-centric analysis. We explore TTS-based speech synthesis and LLM-based text generation to address data scarcity. By disentangling the effects of language confusion and accent bias, our findings demonstrate that syntactic and lexical characteristics exert a more significant influence on CS-ASR performance than speaker accents.
  • We introduce a simplified Equivalence Constraint Theory (SECT) prompting strategy that bridges formal linguistic theory with LLM capabilities. This method outperforms existing prompting methods in linguistic quality and downstream ASR performance.
  • By comparing model- and data-centric approaches, we summarize practical strategies for different code-switching scenarios, offering insight for advancing CS-ASR research and applications.
@@ II-A Unified model and auxiliary tasks
However, the above approach rarely makes explicit use of language information. To mitigate this limitation, recent studies have explored enriching the model via multi-task learning with auxiliary language-aware tasks. As shown in Figure 2(a), these tasks are closely related to language information, such as LD or LID [64, 51, 27]. These tasks are typically performed at the frame or token level to achieve fine-grained language change detection. The CS-ASR model is then optimized with a multi-task objective, formulated as the weighted sum of the ASR loss and the auxiliary language-aware loss as
@@ II-B Language-specific Modules
Another important line of research focuses on incorporating language-specific processing within CS-ASR models, as presented in Figure 2(b). Language-specific encoders process the acoustic feature vectors $\mathbf{X}$ separately into hidden representations $\mathbf{H}=(\mathbf{h}_{t}\in\mathbb{R}^{D}\mid t=1,\ldots,T^{\prime})$ before performing a weighted sum of them. The weights $\mathbf{A}^{mat/emb}=(a_{t}\in[0,1]\mid t=1,\ldots,T)$ are frame-level language posteriors computed by a built-in LID module. These are computed via
@@
The auxiliary language-aware task is typically assigned a relatively small weight (i.e., high $\alpha$ value) compared to the ASR task during training, indicating its minor role in guiding the optimization. By contrast, language-specific modules contribute symmetrically across languages. Each module processes features in the same manner [29, 48], regardless of the relative proportions of matrix and embedded languages in the training data. In the training stage, language-specific modules may require pre-training on constituent languages to learn language-discriminative information before being applied to code-switching data, whereas multi-task learning can be performed solely on code-switching data [51, 27, 28]. During inference, the auxiliary language branch is typically inactive, the decoding process is thus identical to that of a unified model. In contrast, language-specific modules require two passes through the encoder or decoder, introducing additional computational overhead and higher decoding latency. For example, the bi-encoder method exhibits real-time factors (RTF) and latency more than twice those of a model with a single encoder. These differences highlight that auxiliary tasks can enhance CS-ASR with minimal overhead, while language-specific modules deliver more explicit language discrimination at the expense of efficiency.
@@ III-A Data Scarcity and Monolingual Utilization
To address this challenge, existing research has leveraged resources from the constituent languages and explored methods for synthesizing code-switching text and speech. Several studies have shown that monolingual data, particularly from the matrix language, can support the development of CS-ASR systems [29, 58, 34, 57, 68, 53]. For example, [34, 57, 68] investigated building CS-ASR models using only monolingual data. These approaches typically employ language-specific connectionist temporal classification (CTC) [13, 57]. Here, decoding is first performed independently for each language and then combined through joint decoding. Tokens in the matrix-language sequence are then substituted with the corresponding embedded-language tokens when appropriate. However, these works also indicate that subsequent fine-tuning on code-switching data provides substantial performance gains. Moreover, an excessive reliance on monolingual data can result in training imbalance and degrade performance on code-switching speech. Therefore, augmenting code-switching data is crucial for effectively addressing this scarcity.
@@ III-B Code-switching speech synthesis
Figure 3: The CosyVoice 2 model used for TTS, only the LLM module within CosyVoice 2 is updated during fine-tuning.
@@
To analyze the effects of text style and accent bias on CS-ASR, we simulate diverse code-switching speech–text pairs via TTS. The synthetic speech is generated using varying combinations of target and reference text and speech domains. It thus inherits the acoustic characteristics and speaker identity from the reference (i.e., prompt) speech while adhering to the textual pattern of the target text. This design allows us to isolate and evaluate the individual and combined effects of stylistic and accentual variation on ASR robustness.
@@ III-C Code-switching text generation
CS-ASR is inherently a cross-lingual task, as it involves dynamic alternation between two or more languages within a single utterance or conversation. A natural solution to this challenge is to develop a multilingual ASR model equipped with text translation capabilities across language pairs. However, such an approach typically demands large-scale supervised multilingual and parallel data. However, such data are often unavailable, particularly for spontaneous speech commonly observed in code-switching scenarios. Consequently, the applicability of translation-augmented multilingual ASR models remains limited in real-world settings due to severe data scarcity.
An alternative strategy focuses on directly optimizing ASR models with code-switching data. However, this approach is often limited by data scarcity. Although TTS can alleviate the shortage of speech resources, generating diverse and natural code-switching text remains a major challenge. This difficulty arises from the wide variability in code-switching patterns. These patterns may involve arbitrary combinations of syntactic structures, lexical substitutions, and language change points. Consequently, synthesizing high-quality code-switching text is crucial, as it forms the basis for constructing paired speech–text data via TTS [17, 56].
@@
LLMs present a promising solution for this task due to their inherent multilingual and translation capabilities. However, unconstrained generation may lead to unnatural or linguistically implausible switches [17]. To address this, linguistic constraints such as the ECT can be incorporated to guide generation [39, 41, 7, 20, 45]. ECT constrains code-switching to occur only at points where the grammatical structures of both languages align. When implemented in prompt-based generation, ECT can be simplified since LLMs already internalize linguistic regularities to maintain syntactic compatibility and naturalness.
@@
As depicted in Figure 4, monolingual sentences are translated from a high-resource matrix language (Mandarin) into their English-Mandarin code-switching counterparts via the prompt-guided LLM. This process is governed by SECT, which is formulated as three rules. The first rule specifies the matrix language and prevents the LLM from making substantial modifications to the original sentence. The second rule introduces a constraint to control the degree of code mixing. This constraint can be adjusted depending on the desired characteristics of the target code-switching text. Finally, the third rule ensures that the generated code-switching sentences adhere to the syntactic structure of the matrix language while embedding content words from the embedded language at linguistically appropriate switch points. To further guide the model, we adopt an in-context learning strategy by providing three illustrative examples within the prompt. Each example consists of an input and output, being a monolingual sentence and its code-switching counterpart.
Compared with existing LLM-based approaches for code-switching text generation [20, 56], SECT offers a more convenient solution. This is achieved by exploiting the LLM’s intrinsic knowledge and eliminating the need for feeding parallel matrix–embedded language text into each prompt. By leveraging the generative capability of LLMs under SECT constraints, the proposed method synthesizes high-quality code-switching text directly from monolingual data. This generated text can then be used for downstream ASR training.
@@ IV-A Datasets
The two code-switching datasets differ primarily in accent and syntactic structure. Since the ASRU dataset is recorded in mainland China, data within this corpus is characterized by a Chinese accent and contains Mandarin-centric text, with a grammatical structure primarily framed in Mandarin and embedded English segments. Sentences within the ASRU training and development sets, on average, contain 1.6 English words and 8.6 Chinese characters. As opposed to the ASRU data, the SEAME dataset is recorded in Singapore and Malaysia and features speakers with South-East Asian accents. We also compute the code mixing index (CMI) as defined in [12], to measure the degree of language mixing of code-switching text data. The CMI values of the ASRU and SEAME training sets are 17.0 and 13.6, respectively, while the CMI of SEAME training data increases to 24.2 when monolingual utterances are excluded. Therefore, SEAME data contains more frequent code-switching than the ASRU dataset, largely due to the bilingual education systems and language policies in these regions [10]. The higher frequency and fluidity of language switching suggest that SEAME poses greater challenges for code-switching ASR, particularly as the matrix and embedded languages can alternate across sentences [3, 1].